A university bookstore recently sold a wirebound graph-paper notebook for $2.64, and a college-ruled notebook for $3.61. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $168.86. How many of each type were sold?
Answer by rwm(914)

number of wired =w
number of college ruled =c
w+c=50
2.64w+3.61c=168.86
c=38 w=12
38*3.61+12*2.64=168.86